OpenStudy (anonymous):

write the slope intercept form of the equation of the line through the points given.through=(4,2) slope=-1/4.i knw it starts out y=-1/4 b/c y=mx+b

OpenStudy (anonymous):

Good...

OpenStudy (anonymous):

m=1/4 x=4, y=2

OpenStudy (anonymous):

Now plug in the x value (4) for the x and the y value (2) for the y and solve for b

OpenStudy (anonymous):

m=-1/4

OpenStudy (anonymous):

2 = -1/4(4) + b 2 = -1 + b 3 = b

OpenStudy (anonymous):

so you now have m - 1/4 and b = 3 what is your equation?

OpenStudy (anonymous):

did you get the equation

OpenStudy (anonymous):

y=-1/4x+3